Calculate the wavelength(in metres)/associated with an 80g mass moving at 3km/hr
1
Expert's answer
2019-04-03T05:03:00-0400

Wavelength= h / mv = 6.63 x 10 -34 / 80 x 0.278 = 2.98 x 10 - 35